OpenStudy (anonymous):

Solve 5^(x + 5) = 9^x

OpenStudy (anonymous):

Do you know the values of log3 and log5??

OpenStudy (anonymous):

im not sure what u mean, no

OpenStudy (anonymous):

Here to solve this we have to use Logarithm.. Taking Log to the base 10 both the sides, \[Log(5)^{x+5} = Log(9)^x\] Now using power rule of Log, \[Log(x)^a = a.Logx\] So, \[\large (x+5)Log5 = Log(3)^{2x}\] \[\large (x+5)Log5 = 2x.Log3\] Put the values of Log5 and Log3 and then solve for x..

OpenStudy (anonymous):

\[\large Log5 = 0.6989\] \[\large Log3 = 0.4771\] Put these values and then solve for x..

OpenStudy (anonymous):

ok, i got it, thanks :)

OpenStudy (anonymous):

Welcome dear..